OpenStudy (azionne):

4(2x -3) + 8(x - 4) = 2(2x + 6)

pooja195 (pooja195):

Start by distributing \[\rm~(4)(2x)+(4)(−3)+(8)(x)+(8)(−4)=(2)(2x)+(2)(6)\]

OpenStudy (azionne):

ok

OpenStudy (azionne):

shoot im stuck

pooja195 (pooja195):

With?

OpenStudy (azionne):

idk where i messed up at ?

pooja195 (pooja195):

Show me the work .

OpenStudy (pawanyadav):

@pooja195 let me write something ,you are so fast

pooja195 (pooja195):

I'm not fast i've been waiting 5 mins for the asker to write their reply. But go ahead you can take over if you wish :)

OpenStudy (sasquat):

u should get that when u distributed, solve from there 8x-12+8x-32=4x+12

pooja195 (pooja195):

It would be a good thing to see the askers work. To see what they know.

OpenStudy (sasquat):

o ok

OpenStudy (sasquat):

srry

pooja195 (pooja195):

It's alright :)

OpenStudy (azionne):

STOP LET ME DO THIS!! im a little slow today

pooja195 (pooja195):

lol

OpenStudy (azionne):

give up im not feeling it today

OpenStudy (sasquat):

bruh c'mon

OpenStudy (azionne):

i give up bruh im sorry been worken all day ill try again tomorrow

pooja195 (pooja195):

:(